On threads of equal lengths of 50 cm and



negligible masses, about the same point



two balls of equal mass were hung



0.3 g each. When the balls are charged



equal amounts of the same



charges are reflected five times



less distance than length



the end. Calculate the charge of one



balls.

1
Expert's answer
2022-03-13T18:46:34-0400

Explanations & Calculations


  • The electrostatic force between the charges according to Coulomb's law when there is a "\small d" separation between them is

"\qquad\qquad
\begin{aligned}
\small F&=\small k\frac{Q\times Q}{d^2}\\
\small F&=\small k.\frac{Q^2}{d^2}\cdots\cdots(1)
\end{aligned}"

  • Now, from the equilibrium of a single ball-thread system if the angular displacement measured from the vertical is "\small \theta" is

"\qquad\qquad
\begin{aligned}
\small \uparrow\\
\small T\cos\theta&=\small mg\\
\to\\
\small T\sin\theta&=\small F\\
\\
\small F&=\small mg.\tan\theta
\end{aligned}"

  • This into (1) yeilds

"\qquad\qquad
\begin{aligned}
\small k.\frac{Q^2}{d^2}&=\small mg.\tan\theta\\
\small Q&=\small \pm \,d\sqrt{\frac{mg\tan\theta}{k}}\cdots(Answer)\\
\\
\small \sin\theta&=\small \frac{d/2}{l}\\
\small \theta&=\small \sin^{-1}(d/2l)
\end{aligned}"

  • Once you know how much the balls are separated you can calculate the angle and then the tangent of it and finally the charge of a ball.